OpenStudy (anonymous):

Two birds sit at the top of two different trees that are both 29.7 feet tall. The distance between the first bird and a bird watcher on the ground is 42.1 feet. What is the angle measure, or angle of depression, between this bird and the bird watcher?

OpenStudy (asnaseer):

Assuming the distance of 42.1 feet between the first bird and the bird watcher on the ground means "horizontal distance", then this is the diagram for this problem:  You can find the angle of depression \(\theta\) as follows:\[\tan(\theta)=\frac{29.7}{42.1}\]therefore:\[\theta=\tan^{-1}(\frac{29.7}{42.1})\]